Gender Gender
Preference Female Male Total

Coke

85 135

220

Pepsi 120 80 200
Neither/Unsure 65 15 80
Total 270 230 500
  1. If an American is selected at random, what is the probability that the individual is male or prefers Pepsi?
    1. 0.43
    2. 0.16
    3. 0.70
    4. 0.35
  2. If a male respondent is asked which soda he prefers, what is the probability that he will say Coke?
    1. 0.84
    2. 0.16
    3. 0.58
    4. 0.35
Answer #1

Answer - 1: P(Male or prefers Pepsi)

= P(Male) + P(Pepsi) - P(Both)

= 230/500 + 200/500 - 80/500

= 350/500

= 0.70

Option C is correct.

Answer - 2: P(Coke | Male)

= P(Coke and Male) / P(Male)

= 135/230

= 0.58

Option C is correct.